# What You See Is Not Always What You Get: An Empirical Study of Code Comprehension by Large Language Models

### v2

@@ 6.2. Contributions and implications
The methodology of this study draws on practices of preceding studies. However, efforts have been made to reduce the amount of qualitative analysis in data gathering, so that results are not hampered by the subjective confounding variables inherent to human analysis. Particular emphasis is placed on not gathering data on model confidence through model self-reporting, instead, using quantitative tools to provide evidence in this regard. This study made use of the ‘logprobs’ property of the Chat Completions API so that conclusions regarding model confidence can be drawn from quantitative data rather than qualitative model self-reporting – it is the author’s hope that subsequent studies will implement similar quantitative analysis procedures regarding confidence.
6.3. Assumptions and limitations
We assume a ‘No’ response means the LLM does not ‘understand’ the code at all
@@ 6.3. Assumptions and limitations
This study was designed with the assumption that a ‘No’ response is a strong indication that the large language model does not understand the code at all. The ‘ideal behaviour’ of the model in the event of an imperceptible prompt injection attack was defined as it being able to parse it as a human does – i.e.: ignore or ‘reverse-engineer’ the perturbed code and parse the ‘cleaned’ version. However, the results of the study – especially the discrepancy between the GPT-3.5 models and the GPT-4o model – show that this projected ‘ideal behaviour’ is of dubious import, and that it may rather be better to create a system which can guarantee no false negatives, rather than only being able to answer a subset of adversarial examples.
Log probabilities are a good indicator of confidence
@@
A secondary objective of this research has been to harden data gathering procedures against confounding variables which are exhibited in extant research. LLM self-evaluation has been used in preceding studies to record LLM confidence scores, but an explicit requirement in the methodology design of this study is to avoid self-reporting to side-step potential confounding variables inherent in relying upon a closed-source system to evaluate itself. However, it may be the case that log probabilities are also not a perfect representation of model confidence.
@@
What the log probabilities represent is the probability a token is chosen to appear in a specific context (log, 2024). For example, imagine a model is prompted with the straightforward question: ‘Is the sky blue?’ and it returns a single-token response ‘Yes’. This token has a log probability which converts to a linear percentage value of 90%. What that log probability tells us is that this hypothetical model had a 90% chance to respond with ‘Yes’, and a 10% chance to respond with anything else. Now, let us imagine that we prompt the model with a more subjective question such as: ‘Is the Maserati Quattroporte a beautiful car?’ and it again returns a single-token response ‘Yes’. The log probability of that response will be lower, as it is more likely that the model responds ‘No’, even if it ultimately still returned the same response as before. We take this difference between the likelihood of an answer to represent the fall in confidence of the model. Whilst this practice is supported by official documentation (log, 2024), independent research disagrees. One such example is a paper by Panickssery et al. (Panickssery et al., 2024) which indicates that LLMs recognize and favour their own generations even when making evaluations on log probabilities of response tokens. This is why the overall assumption is included here, as a large part of the validity of the gathered data rests on the assumption that log probabilities are an accurate representation of model confidence.
The confidence scores of wrong answers are recorded as -100
@@
In the design of this study, in the event the response is different from what we know the base truth to be, the confidence score of that response is recorded as -100. While it becomes easier to draw conclusions on overall correctness, and confidence scores in correct responses, it normalizes the data in a way which makes it difficult to draw conclusions on the causes of incorrect responses. This design was made with the assumptions outlined in the previous section in mind, and new insights into the veracity of those assumptions may affect the validity of the design of this study.
Heavy limitations placed on responses
@@
The design of this study limits responses to two possible one-token-long responses: either ‘Yes’ or ‘No’. This limitation was enacted so that there would be fewer confounding variables at play when performing log probability confidence evaluation. Preliminary runs of this study design that did not limit token length of response found very high variance in linear probability of responses, as the log probability of each token in the response had to be averaged. This meant that no useful data was gathered with regards to model confidence by using log probabilities, and it would have necessitated human analysis of qualitative results. Since quantitative analysis was greatly preferred to cut down on confounding variables, the decision was made to limit the response length and response options.
6.4. Future work
@@
The greatest limitation faced by this study was finding a way to quantitatively assess model confidence. It is suggested to use novel and effective practices or tools in the future that can better guarantee the validity of confidence scores to produce data from which concrete conclusions can be drawn.
@@
Future research should also delve into how we can create LLM-based systems that can generate a correct response even if the prompt exhibits certain level of perturbations. This is to bridge the gap between what the user experiences and expect, and what the model truly ‘understands’. Guardrails have been shown to negate false positives entirely, which is a marked improvement upon previous generations of LLMs. However, a truly sophisticated natural language processor should be able to follow the spirit of a prompt, rather than the letter of it.
7. Conclusion
@@
The findings in this paper identify the vulnerabilities in recent LLMs for imperceptible coding character attack. With the black-box imperceptible coding character attack framework, we have introduced four distinct attack strategies that effectively manipulate the LLMs’ behaviour in coding related tasks. We have conducted large-scale experiments to demonstrate the susceptibility of LLMs against code snippet perturbations. Furthermore, we have observed that perturbation budget had a strong negative correlation to performance outcomes for code analysis and comprehension task. Additionally, concerning the latest LLMs solution, there is a strong negative correlation between the presence of perturbation and performance outcomes, but no valid correlational relationship between perturbation budget and performance outcomes. We claim that the current devised solutions of imperceptible coding character attacks have presented significant impacts for the coding tasks. We anticipate future research will focus on bridging such issue to enhance the trust of LLMs, in particular for code analysis and comprehension tasks.

### v3

@@ V-B Contributions and implications
The methodology of this study draws on practices of preceding studies. However, efforts have been made to reduce the amount of qualitative analysis in data gathering, so that results are not hampered by the subjective confounding variables inherent to human analysis. Particular emphasis is placed on not gathering data on model confidence through model self-reporting, instead, using quantitative tools to provide evidence in this regard. This study made use of the “logprobs” property of the Chat Completions API so that conclusions regarding model confidence can be drawn from quantitative data rather than qualitative model self-reporting – it is the author’s hope that subsequent studies will implement similar quantitative analysis procedures regarding confidence.
V-C Assumptions and limitations
@@
We assume a “No” response means the LLM does not “understand” the code at all. This study was designed with the assumption that a “No” response is a strong indication that the large language model does not understand the code at all. The “ideal behaviour” of the model in the event of an imperceptible prompt injection attack was defined as it being able to parse it as a human does, i.e. ignore or “reverse-engineer” the perturbed code and parse the cleaned version. However, the results of the study, especially the discrepancy between the GPT-3.5 models and the GPT-4o model, show that this projected “ideal behaviour” is of dubious import, and that it may rather be better to create a system which can guarantee no false negatives, rather than only being able to answer a subset of adversarial examples.
@@
Log probabilities are a good indicator of confidence. A secondary objective of this research has been to harden data gathering procedures against confounding variables which are exhibited in extant research. LLM self-evaluation has been used in preceding studies to record LLM confidence scores, but an explicit requirement in the methodology design of this study is to avoid self-reporting to side-step potential confounding variables inherent in relying upon a closed-source system to evaluate itself. However, it may be the case that log probabilities are also not a perfect representation of model confidence.
@@
What the log probabilities represent is the probability a token is chosen to appear in a specific context [logprobs]. For example, imagine a model is prompted with the straightforward question: “Is the sky blue?” and it returns a single-token response “Yes”. This token has a log probability which converts to a linear percentage value of 90%. What that log probability tells us is that this hypothetical model had a 90% chance to respond with “Yes”, and a 10% chance to respond with anything else. Now, let us imagine that we prompt the model with a more subjective question such as: “Is the Maserati Quattroporte a beautiful car?” and it again returns a single-token response “Yes”. The log probability of that response will be lower, as it is more likely that the model responds “No”, even if it ultimately still returned the same response as before. We take this difference between the likelihood of an answer to represent the fall in confidence of the model. Whilst this practice is supported by official documentation [logprobs], independent research disagrees. One such example is a paper by Panickssery et al. [Panickssery24] which indicates that LLMs recognise and favour their own generations even when making evaluations on log probabilities of response tokens. This is why the overall assumption is included here, as a large part of the validity of the gathered data rests on the assumption that log probabilities are an accurate representation of model confidence.
@@
The confidence scores of wrong answers are recorded as “-100”. In the design of this study, in the event the response is different from what we know the base truth to be, the confidence score of that response is recorded as “-100”. While it becomes easier to draw conclusions on overall correctness, and confidence scores in correct responses, it normalises the data in a way which makes it difficult to draw conclusions on the causes of incorrect responses. This design was made with the assumptions outlined in the previous section in mind, and new insights into the veracity of those assumptions may affect the validity of the design of this study.
@@
Heavy limitations placed on responses. The design of this study limits responses to two possible one-token-long responses: either “Yes” or “No”. This limitation was enacted so that there would be fewer confounding variables at play when performing log probability confidence evaluation. Preliminary runs of this study design that did not limit token length of response found very high variance in linear probability of responses, as the log probability of each token in the response had to be averaged. This meant that no useful data was gathered with regards to model confidence by using log probabilities, and it would have necessitated human analysis of qualitative results. Since quantitative analysis was greatly preferred to cut down on confounding variables, the decision was made to limit the response length and response options.
V-D Potential Threats and Defense Strategies
While imperceptible character-based perturbations may sometimes prevent code from compiling, our goal is not to evaluate the correctness of the code snippets. Furthermore, it does not eliminate their security relevance for the investigation of LLMs’ capability, especially in workflows where LLMs are used to review or validate code before it is ever executed. We highlight two key threat scenarios and propose defense strategies that address them at various stages of the LLM-assisted code analysis pipeline.
Threat 1: Post-copy vulnerability propagation. In real-world workflows, code is often copied between platforms (e.g., from a website to an IDE, or from a chat to a version control system). During this process, certain imperceptible characters may be stripped or normalised, resulting in a clean-looking version of the code that compiles and runs, but one that was previously reviewed and falsely approved by an LLM. This creates a false sense of safety, as neither the LLM nor the human reviewer recognised the original obfuscation.
Threat 2: User confusion and resource waste. When perturbed code does survive intact, developers may face unclear compilation errors or behavioural inconsistencies. These issues can lead to time-consuming debugging, misdiagnosed problems, or abandonment of useful code. In high-pressure environments, such disruptions erode trust in AI code review tools and increase reliance on manual inspection.
To mitigate these risks, we recommend the following defense strategies:
1. Pre-processing and normalization: Implement input sanitisation steps such as Unicode normalisation, removal of zero-width characters, and visual similarity checks (e.g., homoglyph detection). These measures reduce the likelihood of LLMs encountering malformed or deceptive input.
2. Multi-pass verification: Avoid relying solely on a single LLM judgment. Instead, aggregate results from multiple models or rerun the same model under varied conditions. Only proceed with critical decisions when consensus or confidence thresholds are met.
3. Explainability enforcement: Require LLMs to justify their reasoning when labelling code as safe or unsafe. Structured explanations can help detect hallucinated judgments, expose oversights, and provide transparency for both humans and downstream tools.
These strategies are not intended to guarantee complete robustness. Rather, they offer a layered defense framework tailored to the unique risks posed by imperceptible code perturbations in LLM-centered pipelines. Addressing these subtleties is essential for building trustworthy, secure AI-assisted development environments.
V-E Future work
@@
The primary limitation of this study lies in finding a way to quantitatively assess model confidence. It is suggested to use novel and effective practices or tools in the future that can better guarantee the validity of confidence scores to produce data from which concrete conclusions can be drawn.
@@
Future research should also delve into how we can create LLM-based systems that can generate a correct response even if the prompt exhibits certain level of perturbations. This is to bridge the gap between what the user experiences and expect, and what the model truly “understands”. Guardrails have been shown to negate false positives entirely, which is a marked improvement upon previous generations of LLMs. However, a truly sophisticated natural language processor should be able to follow the spirit of a prompt, rather than the letter of it.
VI Conclusion
@@
The findings in this paper identify the vulnerabilities in recent LLMs for imperceptible coding character attack. With the black-box imperceptible coding character attack framework, we have introduced four distinct attack strategies that effectively manipulate the LLMs’ behaviour in coding related tasks. Our primary goal is to assess whether LLMs can maintain alignment with the original code intent with minimal visual perturbations, which perturbation may not result in functionally executable code snippets. We have conducted large-scale experiments to demonstrate the susceptibility of LLMs against code snippet perturbations. Furthermore, we have observed that perturbation budget had a strong negative correlation to performance outcomes for code analysis and comprehension task. Additionally, concerning the latest LLMs solution, there is a strong negative correlation between the presence of perturbation and performance outcomes, but no valid correlational relationship between perturbation budget and performance outcomes. We claim that the current devised solutions of imperceptible coding character attacks have presented significant impacts for the coding tasks. We anticipate future research will focus on bridging such issue to enhance the trust of LLMs, in particular for code analysis and comprehension tasks.